# An entanglement witness and an effective criterion for bipartite states of different dimensions

### v4

@@ Abstract
We present a simple approach for generation of a diverse set of positive maps between spaces of different dimensions. The proposed method enables the construction of Entanglement Witnesses tailored for systems in $d_{1}\times d_{2}$ dimensions. We then show how application of this entanglement witness to a general density matrix leads to a simple calculable criterion in closed form. As a support for the power of this criterion, we apply it to a variety of (one-parameter) PPT entangled states and show that the ranges of the parameters where these states are entangled are much larger than previously reported.
@@ Introduction
Entanglement stands out as one of the most remarkable distinctions between the quantum and classical realms, playing a pivotal role in numerous quantum protocols, for review papers see [1, 2, 3]. As a result, the creation, manipulation, application, and identification of entanglement have profound significance. The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [4]. Hence, the pursuit of viable techniques for detecting entangled states takes on paramount importance.
Entangled states are referred to as states that are not separable, and separable states refer to states that can be written as follows[5]:
@@
where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$. One strategy for identifying entanglement involves investigating whether a state can be expressed in the aforementioned form. This, however, poses a formidable challenge. As a result, there is a pressing need for more effective methods for the detection of entanglement. Over time, diverse techniques have emerged for this purpose. The Peres criterion is considered one of the primary methods in this regard[6]: It’s relatively straightforward to see that the partial transpose of all separable states is positive. This is seen simply by noting that for any state of the form (1), one has
@@
where $T$ is the transpose, hence the name partial transpose. Therefore, if the partial transpose of a quantum state is not positive, it must be entangled. However, this criterion alone cannot identify all types of quantum entangled states because there are entangled states known as PPT, whose partial transpose are positive. In fact separable states are a subset of PPT states and only when $d_{1}d_{2}\leq 6$, these two sets are equal. Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems. The effectiveness of the Peres criterion (2) hinges on a basic property of the transpose map $T$ which is a positive and yet not completely positive map. Therefore it was natural to search for more general types of positive maps lacking the property of complete positivity and thus constructing more general Entanglement Witnesses than the simple partial transpose. In fact, it has been known that any entanglement witness $W$ can be written in the form
$W(M)=(I\otimes M)|\phi^{+}\rangle\langle\phi^{+}|,$
where $M$ is a positive map and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[7]. In other words, $W(M)$ is the Choi-matrix of the map $M$, which due to its lack of complete positivity can have negative eigenvalue. Such a witness has the property
@@
where $S$ is the set of all separable states. Alternatively, one can think of $W$ as a block-positive operator, i.e.
@@
which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$. Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps. Based on this equivalence, entanglement witnesses were introduced in many works, see [7, 8, 9, 10] and references therein. For more recent works, see [11, 12, 13]. In practical situations, where the two particles are far apart in remote laboratories, application of any entanglement witnesses found so far, requires a common reference frame between the two parties holding the particles. It was recently shown in [14] that by a process called incomplete teleportation, all these entangled states are still able to detect entanglement of remote parties, even in the absence of a common frame of reference.
In contrast with completely positive maps which due to the Kraus theorem can be easily constructed, construction of positive maps is not a straightforward task, and appropriate methods need to be proposed for this purpose. In addition to the finding new classes of witnesses, the relation of EW’s and positive maps, have led to many new insights on both [8]. For example, a basic distinction refers to decomposable witnesses which are of the form
$W=P+Q^{\Gamma},$
where $P,Q$ are positive operators and $\Gamma$ is the operator of partial transposition. Any witness that cannot be written in the above form is called indecomposable. It can be proved that decomposable witnesses cannot identify PPT entangled states. In this sense these are somewhat weaker than indecomposable witnesses. One can also characterize witnesses by their optimality and extremality [8], a subject which we will not deal with in this work. The line that we follow is inspired by a number of recent works [15, 16, 17, 18, 19, 20] in which specific types of positive maps have been defined and the corresponding witnesses have been constructed.
It is worth noting that in almost all the mentioned articles, the entanglement witnesses constructed have been designed for two systems of equal dimensions. In this paper, we propose a simple method for identification of entanglement in systems with dimensions of $d_{1}\times d_{2}$. This dimension-specific applicability enhances the versatility of the method, making it a valuable tool in different practical scenarios. Often, due to practical constraints in laboratory setups, certain measurements may be more accessible or accurate than others. Our method elegantly accommodates this by allowing for the construction of entanglement witnesses based solely on those restricted measurements which are feasible in the given experimental context(See examples).
The structure of this paper is as follows: In Section 1, we examine the main idea of this article, which is to construct a new class of positive maps. In Section 2, an alternative method is utilized to obtain the same results that presented in Section 1. In section 3 we provide a simple criterion built with our Entanglement Witness which easily detects entangled states. In Section 4, we present several examples which shows the power of our criterion in detecting PPT entangled states. We end up the paper with a conclusion.
1 Construction of positive maps for arbitrary different dimensions
Consider the Hilbert space $H_{d_{1}}$ of dimension $d_{1}$ and let a pure state $X\in D(H_{d_{1}})$, (the space of states on $H_{d_{1}}$) be written as
$X=\frac{1}{d_{1}}I_{d_{1}}+\sum_{i=1}^{d_{1}^{2}-1}x_{i}\Gamma_{i},$
where $\{\Gamma_{i},\ \ \ i=1\cdots d_{1}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{1}})$, i.e.
$\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij},\ \ \ \operatorname{Tr}\Gamma_{i}=0.$
Purity of the state $X$ (i.e. $\operatorname{Tr}X^{2}=1$) and the above normalization conditions restricts the norm of the vector ${\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1})$ to
${\bf x}\cdot{\bf x}=\frac{d_{1}-1}{d_{1}},$ (3)
where ${\bf x}\cdot{\bf x}:=\sum_{i=1}^{d_{1}^{2}-1}x_{i}^{2}$.
Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be unital and trace-preserving map which acts as follows:
$\Phi(\frac{1}{d_{1}}I_{d_{1}})=\frac{1}{d_{2}}I_{d_{2}}\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\frac{1}{d_{2}}\sum_{k=1}^{d_{2}^{2}-1}\Lambda_{ik}\Omega_{k}.$ (4)
In fact, $\Lambda_{ik}$ are real numbers representing the characterization of the map $\Phi$. Here, our definition remains consistent with the trace-preserving property of the map, in contrast to the definition presented in [18], which is grounded in the property of such a map that takes an identity matrix in $D(H_{d_{1}})$ to a identity matrix in $D(H_{d_{2}})$. The factor $\frac{1}{d_{2}}$ in the second equation is written for later convenience, and $\{\Omega_{k},\ \ \ k=1\cdots d_{2}^{2}-1\}$ are a set of orthonormal traceless Hermitian operators $\in L(H_{d_{2}})$, i.e.
$\operatorname{Tr}(\Omega_{k}\Omega_{l})=\delta_{kl},\ \ \ \operatorname{Tr}\Omega_{k}=0.$
Then $\Phi$ maps $X$ to
$\Phi(X)=\frac{1}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$ (5)
Remark 1
Up to now we have assumed that $X$ is a pure state. This can be extended to arbitrary operatores $X\in L(H_{d_{1}})$ as follows:
$\Phi(X)=\frac{\operatorname{Tr}(X)}{d_{2}}I_{d_{2}}+\frac{1}{d_{2}}\sum_{i=1}^{d_{1}^{2}-1}\sum_{k=1}^{d_{2}^{2}-1}x_{i}\Lambda_{ik}\Omega_{k}.$
To ensure positivity of the map $\Phi$, we now use Mehta’s Lemma [21] which we state below:
Mehta’s Lamma (1989): Let $A$ be a Hermitian matrix of dimension $D$, if
@@ 1 Construction of positive maps for arbitrary different dimensions
In view of linearity of the map $\Phi$, to prove positivity of the map $\Phi$, it is sufficient to prove that $\Phi(X)$ is positive for any pure state $X$. That is the map $\Phi$ is positive if the following condition holds
$\operatorname{Tr}(\Phi(X)^{2})\leq\frac{(\operatorname{Tr}\Phi(X))^{2}}{d_{2}-1}=\frac{1}{d_{2}-1}.$ (7)
To this end, we calculate $\operatorname{Tr}(\Phi(X)^{2})$, which in view of orthonormality condition (4), turns out to be
$\operatorname{Tr}(\Phi(X)^{2})=\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x},$
where the matrix $\Lambda$ is composed of $\Lambda_{ij}$ elements. In view of (6), we demand that
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}{\bf x}^{T}{\Lambda}\Lambda^{T}{\bf x}\leq\frac{1}{d_{2}-1},\hskip 28.45274pt\forall{\bf x}.$
We can now replace this with an even stronger inequality, namely
$\frac{1}{d_{2}}+\frac{1}{d_{2}^{2}}\lambda_{max}(\Lambda{\Lambda^{T}}){\bf x}\cdot{\bf x}\leq\frac{1}{d_{2}-1},$ (8)
where $\lambda_{max}:=\lambda_{max}(\Lambda\Lambda^{T})$, is the largest eigenvalue of the matrix $\Lambda\Lambda^{T}$, or equivalently $\lambda_{max}$ is the square of the largest singular value of the matrix $\Lambda^{T}$. Inserting (8) in (7) and using (3), yields the following constraint on $\lambda_{max}$
$\lambda_{max}\leq\frac{d_{1}d_{2}}{(d_{1}-1)(d_{2}-1)}.$ (9)
The EW corresponding to the map (5) is given by [7]

### v5

# Positive maps and Entanglement Witnesses in different dimensions

@@ Abstract
We present a continuous, multiparameter family of positive maps between spaces of differing dimensions. This framework facilitates the construction of Entanglement Witnesses (EWs) specifically designed for systems in $d_{1}\times d_{2}$ dimensions. We derive a simple, closed-form criterion for detecting entanglement in general density matrices based on these witnesses. To demonstrate the effectiveness of this criterion, we apply it to a range of Positive Partial Transpose (PPT) entangled states, revealing that the parameter regions where these states exhibit entanglement are larger than previously reported. Furthermore, we prove that non-unital EWs, corresponding to non-unital maps, are not more powerful than unital EWs, thus supporting the focus on unital positive maps in recent studies. Our method complements existing approaches to separability criteria for density matrices in different dimensions.
@@ Introduction
Any conceivable operation on a quantum system, be it a measurement or noisy evolution, can be represented by a Completely Positive Trace-preserving (CPT) map or a quantum channel for short. A map of this kind not only maps density matrices to density matrices, but all of its extensions also have this property. The name of completely positive map comes from this last property. At first sight it seems, due to this stringent completely positive property, that construction of CPT maps is much more difficult than that of simply positive maps. However it has turned out that it is quite the contrary. CPT maps can readily be constructed, among other methods, by resorting to their Kraus representation [1]. In fact any such map from a system $A$ to another system $B$ (possibly with different dimension can be written as
${\cal E}(\rho)=\sum_{\alpha=1}^{K}A_{\alpha}\rho A_{\alpha}^{\dagger}$ (1)
where $A_{\alpha}:H_{A}\longrightarrow H_{B}$ are arbitrary linear operators whose only condition is the trace-preserving property, $\sum_{\alpha=1}A_{\alpha}^{\dagger}A_{\alpha}=I_{A}$, a condition which can be easily fullfiled, i.e. one simply chooses arbitrary operators $A_{\alpha},\ \ \alpha=1,\cdots K-1$ and by multiplying them with suitable scale factors ensures that $\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}\leq I_{A}$ and defines the last Kraus operator simply as $A_{K}:=\sqrt{I-\sum_{\alpha=1}^{K-1}A_{\alpha}^{\dagger}A_{\alpha}}$. While construction of instances of CPT maps is not challenging, classification of these maps, specially in higher dimensions, and deciding whether or not a given map is CPT is generally difficult. On the other hand construction of instances of maps which are only positive, but not completely positive, is not straightforward at all and as expected, their classification has remained an open problem despite decades of efforts by physicists and mathematicians. The need for construction of instances of positive maps arises from their role in entanglement theory [2, 3, 4].
The identification of entangled states, in particular, presents a formidable challenge, underscored by its classification as an NP-Hard problem in mathematical terms [5]. Hence, the pursuit of viable techniques for detecting entangled states takes on paramount importance. Entangled states are referred to as states that are not separable, and separable states refer to states that can be written as follows[6]:
@@
where $0\leq p_{i}\leq 1$ and $\sum p_{i}=1$. It is not easy at all to determine whether or not a given state can be expressed in the aforementioned form. As a result, there is a pressing need for more effective methods for the detection of entanglement. Over time, diverse techniques have emerged for this purpose. The Peres criterion is considered one of the primary methods in this regard[7], i.e. a state which is separable satisifes
@@
where $T$ is the transpose, hence the name partial transpose. Therefore, if the partial transpose of a quantum state is not positive, it must be entangled. However, this criterion alone cannot identify all types of quantum entangled states because there are entangled states known as PPT, whose partial transpose are positive. In fact separable states are a subset of PPT states and only when $d_{1}d_{2}\leq 6$, these two sets are equal. Here $d_{1}$ and $d_{2}$ are the dimensions of the two quantum systems. In higher dimensions one should search for other criteria of entanglement, many of which have been reviewed in [8]. The most important ones are called Entanglement Witnesses (EW), see [9] for a review. An EW refers to a Hermitian operator $W$, which has negative eigenvalues but has the following property:
@@
where $S$ is the set of all separable states. Therefore, a negative expectation value for $W$ on a state $\rho$, then means that $\rho$ is entangled. Alternatively, one can think of $W$ as a block-positive operator, i.e.
@@
which has negative eigenvalue when acting on the whole Hilbert space $H_{1}\otimes H_{2}$. Here comes the important link between entanlement witnesses and positive maps. In fact, it has been known that any entanglement witness $W$ can be written in the form
$W=(I\otimes\Phi)|\phi^{+}\rangle\langle\phi^{+}|,$
where $\Phi$ is a positive (but not completely positive) maps, and $|\phi^{+}\rangle=\frac{1}{\sqrt{d}}\sum_{\mu=0}^{d-1}|\mu\mu\rangle$, is a maximally entangled state[10]. In other words, $W$ is the Choi-matrix of the map $\Phi$, which due to its lack of complete positivity can have negative eigenvalue. Thus the problem of constructing entanglement witnesses becomes equivalent to constructing positive but not completely positive maps. This is an important and known result which connects the long-standing problem of classification of positive maps in pure mathematics to the new problem of detection of entanglement quantum information theory. Based on this equivalence, entanglement witnesses were introduced in many works, see [10, 9, 11, 8] and references therein. For more recent works in this direction, see [12, 13, 14], where the authors generally start from a restricted class of positive maps, mostly between spaces of equal dimensions and construct the corresponding EW’s. These maps are always uniital and their corresponding affine transformation is a rotation performed by an orthogonal matrix.
In this work, we construct a large class of positive maps which firstly act between spaces of different dimensions $d_{1}\times d_{2}$. Besides detecting entanglement in heterogenuous systems, this makes it possible to witness entanglement in a multi-partite system, for example when $k$ qubits are entnagled to $n-k$ qubits in a many-body system. In particular, we
a) provide sufficient conditions for our maps to be positive, but not completely positive, so that their Choi matrix is an EW, and then
b) provide sufficient conditions for our maps to be completely positive. Both criteria are expressed in terms of the affine transformation corresponding to this map.
c) We then prove that among the EW’s constructed from positive maps, the unital EW’s, that is, those constructed from unital maps are stronger than those corresponding to non-unital maps. Thus we rigorously support the focus on unital positive maps in recent studies [12, 13, 14].
d) We reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses. Thus simply by checking this critria, one can decide whether or not a given state is entangled.
e) By studying several examples, we show that the EW’s generated in this way, can detect entanglement of states for a larger range of parameters than previously found.
f) Finally in an appendix, we restrict ourselves to unital EW’s and show how by using the concept of block-positivity, we can directly construct the EW’s derived from the positive maps.
The structure of this paper is as follows: In Section 1, we set up our notation. Then in section 2, we state Mehta’s lemma and its proof which is the basis of all our construction. In section 3, we use this lemma [15] to construct general positive maps between spaces of different dimensions. By general maps, we mean maps which are not necessarily unital. This part extends the previous works which have constructed EW’s along similar lines [16, 17, 18]. In section 4 we reduce the previous considerations to a simple criterion which solely depends on the density matrix itself and does not require any type of entanglement witnesses. We also show in this section that in each family of maps, the unital ones lead to the strongest type of witnesses. In Section 5, we present several examples which shows the power of our criterion in detecting PPT entangled states. We end up the paper with a conclusion. In an appendix, section 8, an alternative method based on the concept of block-positiviy is used to obtain the same results derived in section 3.
1 Notations and Conventions
Let $H_{d}$ be the Hilbert space of dimension $d$ and $L(H_{d})$ the space of linear operators acting on $H_{d}$. The half space of positive operators is denoted by $L(H_{d})^{+}$. We will be dealing with maps $\Phi:L^{+}(H_{d_{1}})\longrightarrow L^{+}(H_{d_{2}})$. The Hilbert-Schmidt norm of a real matrix $A$ is denoted by $\|A\|=\operatorname{Tr}(A^{T}A)^{1/2}$. By $S\in M_{d_{1}\times d_{2}}$, we mean a rectangular matrix of dimension $d_{1}\times d_{2}$, by $M\in O(d)$, we mean an orthogonal square matrix of dimension $d$. Let $A$ be a rectangular matrix. Then by $a_{max}$ we mean the largest eigenvalue of the positive symmetric matrix $\sqrt{A^{T}A}$ or equivalently $\sqrt{AA^{T}}$. A basis of orthonormal Hermitian and traceless matrices acting on $H_{d_{1}}$ is denoted by $\{\Gamma_{i},\ i=1\cdots d_{1}^{2}-1\}$ and the same basis of matrices acting on $H_{d_{2}}$ is denoted by $\{\Omega_{\mu},\mu=1,\cdots d_{2}^{2}-1\}$. Therefore we have $\operatorname{Tr}(\Gamma_{i})=\operatorname{Tr}(\Omega_{\mu})=0$ and $\operatorname{Tr}(\Gamma_{i}\Gamma_{j})=\delta_{ij}$ and $\operatorname{Tr}(\Omega_{\mu}\Omega_{\nu})=\delta_{\mu,\nu}.$ Finally we use the convention of summing over repeated indices. In all formulas, $i$ runs from $1$ to $d_{1}^{2}-1$ and $\mu$ from $1$ to $d_{2}^{2}-1$.
The state of a quantum system is denoted by $\rho\in D(H_{d})$, where $D(H_{d})\subset L(H_{d})^{+}$ is the convex subset of positive operators with unit trace. Let $X$ be a positive operator in $L^{+}(H_{d_{1}})$ which can be expanded as
$X={x_{0}}I_{d_{1}}+x_{i}\Gamma_{i}=x_{0}I_{d_{1}}+{\bf x}\cdot\bm{\Gamma},$ (4)
where $(x_{0},{\bf x}=(x_{1},\cdots x_{d_{1}^{2}-1}))$ are real parameters. For any positive operator $X$, it necessarily holds that $\operatorname{Tr}(X^{2})\leq\operatorname{Tr}(X)^{2}$ which leads to
${\bf x}\cdot{\bf x}\leq x_{0}^{2}d_{1}(d_{1}-1).$ (5)
where ${\bf x}\cdot{\bf x}:=x_{i}x_{i}$. If the operator $X$ is pure, that is, if $X=|\psi\rangle\langle\psi|$ for a non-normalized vector $|\psi\rangle$, then the equality sign holds in (5).
Any operator $Y$ in $L(H_{d_{2}})$ is written as
$Y={y_{0}}I_{d_{2}}+y_{\mu}\Omega_{\mu}=y_{0}I_{d_{2}}+{\bf y}\cdot\bm{\Omega}.$ (6)
Let $\Phi:D(H_{d_{1}})\longrightarrow D(H_{d_{2}})$ be a map which acts as follows:
$\Phi(I_{d_{1}})=(I_{d_{2}}+t_{\mu}\Omega_{\mu})\ \ \ \ ,\ \ \ \Phi(\Gamma_{i})=\Lambda_{\mu i}\Omega_{\mu},$ (7)
Here $t_{\mu}$ and $\Lambda_{\mu i}$ are real numbers which characterize the map $\Phi$. In the special case where ${\bf t}=0$, the map $\Phi$ is unital. Hereafter we use summation convention for summing over repeated indices. With these conventions, $\Phi$ maps $X$ to
$\Phi(X)={x_{0}}I_{d_{2}}+(x_{0}t_{\mu}+x_{i}\Lambda_{\mu i})\Omega_{\mu}=x_{0}I_{d_{2}}+({\bf t}x_{0}+\Lambda{\bf x})\cdot\Omega.$ (8)
With our convention, the map $\Phi$ corresponds to the following affine tranformation, $\Phi:(x_{0},{\bf x})\longrightarrow(x_{0},{\bf t}x_{0}+\Lambda{\bf x})$ which can be written in matrix form .
$\Phi:\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}\longrightarrow\begin{pmatrix}1&{\bf 0}\\ {\bf t}&\Lambda\end{pmatrix}\begin{pmatrix}x_{0}\\ {\bf x}\end{pmatrix}.$ (9)
where ${\bf t}$ is a $d_{2}$ dimensional vector and $\Lambda\in M_{d_{2}\times d_{1}}$.
2 Mehta’s Lamma
Let $A$ be a Hermitian matrix of dimension $D$, if
@@ 2 Mehta’s Lamma
Note that Mehta’s lemma provides only a sufficient condition. Its converse is not true, e.g. there are positive matrices which do not satisfy (10), as is readily seen from an example of a diagonal matrix $A=diag(1,2,x),\ \forall\ x\geq 6$.
Proof: The proof is taken from refernce [19]. Let $A$ be a Hermitian matrix of size $D$ and let $\alpha={\rm Tr}A/\sqrt{{\rm Tr}A^{2}}$. If $\alpha\geq\sqrt{D-1}$ then $A$ is positive. We calculate the traces in the eigenbasis of $A$. Let $(x_{1},\dots x_{D})$ denote the spectrum of $A$. Assume first that one eigenvalue, say $x_{1}$, is negative. Making use of the right hand side of the standard estimation between the ${l}_{1}$ and ${l}_{2}$–norms (with prefactor $1$) of a $D$–vector, $||A||_{2}\leq||A||_{1}\leq D||A||_{2}$, we infer
${\rm TrA}=\sum_{i=1}^{D}x_{i}<\sum_{i=2}^{D}x_{i}\leq\sqrt{D-1}\Bigl{(}\sum_{i=2}^{D}x_{i}^{2}\Bigr{)}^{1/2}<\sqrt{D-1}\sqrt{{\rm Tr}A^{2}}\ .$ (11)
This implies that $\alpha<\sqrt{D-1}$. Hence if the opposite is true and $\alpha\geq\sqrt{D-1}$ then none of the eigenvalues $x_{i}$ could be negative, so $A\geq 0$. $\square$.
3 Positive maps between spaces of different dimensions
Given a quantum map $\Phi$ characterized by the affine transformation on the generalized Bloch vector ${\bf r}\longrightarrow\Lambda{\bf r}+{\bf t}$, our goal is to find the condition on this affine map which gurantees positivity of $\Phi$. This is stated in the following:
3.1 Sufficient condition for positivity
Proposition: A map $\Phi$, implementing the affine transformatoin (9) is positive if the following condition holds:
$\|{\bf t}\|+\sqrt{d_{1}(d_{1}-1)}\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{2}-1}},$ (12)
where $\lambda_{max}$ is the largest eigenvalue of the positive matrix $\sqrt{\Lambda\Lambda^{T}}.$
Proof: Due to linearity of the map, it suffices to prove the statement only for pure states. For such states (see (5)), we have
$\operatorname{Tr}(\Phi(X))=x_{0}d_{2}$
and
$\operatorname{Tr}(\Phi(X)^{2})=x_{0}^{2}d_{2}+\|{\bf t}x_{0}+\Lambda{\bf x}\|^{2}={x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}.$
In view of the Mehta lemma, the map $\Phi$ is positive if for all $(x_{0},{\bf x})$ subject to the condition (10), the following inequality holds
${x_{0}^{2}}d_{2}+x_{0}^{2}{\bf t}\cdot{\bf t}+2x_{0}{\bf t}\cdot\Lambda{\bf x}+\|\Lambda{\bf x}\|^{2}\leq\frac{x_{0}^{2}d_{2}^{2}}{d_{2}-1}.$ (13)
We demand a slightly stronger condition, (at a price of loosing some of the positive maps). To this end, we note that
$\|\Lambda{\bf x}\|^{2}\leq\lambda_{max}^{2}{\bf x}\cdot{\bf x}=\lambda_{max}^{2}x_{0}^{2}d_{1}(d_{1}-1),$ (14)
where $\lambda_{max}^{2}$ is the largest eigenvalue of the positive matrix $\Lambda^{T}\Lambda$. We also use Cauchy-Schwartz inequality to write
${\bf t}\cdot\Lambda{\bf x}\leq\|{\bf t}\|\ \|\Lambda{\bf x}\|\leq\|{\bf t}\|\lambda_{max}\|{\bf x}\|=\|{\bf t}\|\lambda_{max}{x_{0}}\sqrt{d_{1}(d_{1}-1)},$ (15)
Inserting (14) and (15) in (13) and dividing both sides by $x_{0}^{2}$, we find the following sufficient condition for positivity of the map $\Phi$
$\|{\bf t}\|^{2}+2\lambda_{max}\|{\bf t}\|\sqrt{d_{1}(d_{1}-1)}+\lambda_{max}^{2}d_{1}(d_{1}-1)\leq\frac{d_{2}}{(d_{2}-1)}.$ (16)
Taking the square root of both sides leads to (12) which concludes the proof. $\square$
For unital maps where ${\bf t}=0$, the above condition simplifies to
$\lambda_{max}\leq\sqrt{\frac{d_{2}}{d_{1}(d_{1}-1)(d_{2}-1)}}.$
3.2 Sufficient condition for complete positivity
A map $\Phi$ is completely positive if all its extensions are also positive. We now derive a sufficient condition on the affine map for complete positivity of the corresponding map.
Proposition: A positive map (8), implementing the affine transformatoin (9) is a completely positive map if the following condition holds:
${\bf t}\cdot{\bf t}+d_{1}\operatorname{Tr}(\Lambda\Lambda^{T})\leq\frac{d_{2}}{d_{1}d_{2}-1}.$ (17)
Proof: According to a celebrated theorem of Choi [20], a map is completely positive if its Choi matirx defined as below is positive
$C_{\Phi}$ $=$ $(I_{d_{1}}\otimes\Phi)(|\phi^{+}\rangle\langle\phi^{+}|)=(|i\rangle\langle j|\otimes\Phi(|i\rangle\langle j|)$ (18)
$=$ $|i\rangle\langle j|\otimes\Big{[}\frac{1}{d_{1}}\delta_{ij}(I_{d_{2}}+{\bf t}\cdot\Omega)+\langle j|\Gamma_{l}|i\rangle\Lambda_{\mu l}\Omega_{\mu}\Big{]}$ (19)
$=$ $\frac{1}{d_{1}}I_{d_{1}}\otimes(I_{d_{2}}+{\bf t}\cdot\Omega)+\Lambda_{\mu l}\Gamma_{l}^{T}\otimes\Omega_{\mu}.$ (20)
from which the traces are easily derived as
$\operatorname{Tr}(C_{\Phi})=d_{2},\hskip 28.45274pt\operatorname{Tr}(C_{\Phi}^{2})=\frac{d_{2}+{\bf t}\cdot{\bf t}}{d_{1}}+\operatorname{Tr}(\Lambda\Lambda^{T})$ (21)
Inserting these in Mehta’s Lemma (10), leads to (17) and completes the proof $\square$.
Example: Maps between qubits: As an example consider the qubit maps $\Phi$ with the affine transformation (9) where $\Lambda=\begin{pmatrix}\lambda_{1}&&\\ &\lambda_{2}&\\ &&\lambda_{3}\end{pmatrix}$ and ${\bf t}=(t_{1},t_{2},t_{3})^{T}$. Then the condition for positivity of the map is given from (12) to be
$\sqrt{t_{1}^{2}+t_{2}^{2}+t_{3}^{2}}+\sqrt{2}\lambda_{max}\leq\sqrt{2},$ (22)
and the condition of complete positivity is given from (17) to be
$\sum_{i}(t_{i}^{2}+2\lambda_{i}^{2})\leq\frac{2}{3}.$ (23)
We shoud emphasize that since the Mehta’s condition is a sufficient conditions, the above inequalities determine only a subset of positive and completely positive maps. For example, if we restrict ourselves to unital maps, then the condition (17) is given by
$\lambda_{1}^{2}+\lambda_{2}^{2}+\lambda_{3}^{2}\leq\frac{1}{3},$
which is a sphere of radius $r=\frac{1}{\sqrt{3}}$ and is a subset of the Fujiwara-Algoet tetrahedron
$|\lambda_{1}+\lambda_{2}|\leq 1+\lambda_{3}\ \ \ \ \ |\lambda_{1}-\lambda_{2}|\leq 1-\lambda_{3}.$
A comparison of the volumes of these two sets shows that $30$ percents of the completely positive unital maps for qubits are detected by this simple use of Mehta’s criteria, figure (1).
Figure 1: Example of qubit maps, with eigenvalues of affine transformation given by $\lambda_{1},\lambda_{2}$ and $\lambda_{3}$. Restricting to unital maps, the tethdrahedron shows the complete set of unital maps as derived from the Fujiwara-Algoet condition, the sphere shows those maps which are derived from Mehta’s condition. Its volume is approximately 30 percent of the volume of the tethrahedron.
4 Entanglement Witnesses from positive maps
It is known that there is a one-to-one correspondence between positive map which are not completely positive on the one hand and Entanglement Witnesses on the other. The EW corresponding to the map (8) is given by [10]